A girl has a number of ducks costing $10 each and three times as many chickens costing $6 each. If the total cost of ducks and chickens is $420, find the number of chickens the girl has.
 Answer by haileytucki(390)
       
 A girl has a number of ducks costing $10 each and three times as many chickens costing $6 each. If the total cost of ducks and chickens is $420, find the number of chickens the girl has. x(10)+3x(6)=420
 Multiply x by each term inside the parentheses.
 10x+3x(6)=420
 Multiply 3x by each term inside the parentheses.
 10x+18x=420
 Since 10x and 18x are like terms, add 18x to 10x to get 28x.
 28x=420
 Divide each term in the equation by 28.
 (28x)/(28)=(420)/(28)
 Simplify the left-hand side of the equation by canceling the common factors.
 x=(420)/(28)
 Simplify the right-hand side of the equation by simplifying each term.
 x=15
 Now, if x=15 then 15*10=150 ducks
 | 
  
 | 
 |